# From Contrast to Commonality: Audio Commonality Captioning for Enhanced Audio-Text Cross-modal Understanding in Multimodal LLMs

### v2

@@ 2.1 Construction of the ADC and ACC Dataset
Qwen2-Audio + AC 0.0894 0.0506 0.0295 0.1762 0.5922 0.1529 0.2831 0.4132 0.1005 0.1762
Qwen2-Audio + ADC 0.0262 0.0118 0.0053 0.0798 0.3861 0.0566 0.0754 0.0943 0.0442 0.0798
Qwen2-Audio + ACC (ours) 0.1039 0.0585 0.0338 0.1999 0.5804 0.1534 0.2826 0.4118 0.1032 0.1999
@@ 3.1 Dataset and Metrics
In this study, we investigate the relative impact of direct Audio Captioning (AC) fine-tuning, Audio Difference Captioning (ADC), and Audio Commonality Captioning (ACC) on enhancing the audio-text cross-modal understanding of MLLMs. We adopt the audio captioning task as the primary evaluation benchmark, as it directly reflects the model’s ability to achieve semantic alignment between audio and text before and after fine-tuning. For evaluation, we utilize two widely used datasets—AudioCaps [1] and Clotho [3]. It is worth noting that AudioCaps is partially incorporated into the construction of our fine-tuning datasets for AC, ADC, and ACC through sampling and mixing, which may introduce some overlap between pretraining and evaluation data. Nevertheless, within the context of MLLMs, such strategies remain valuable, as they offer meaningful empirical insights and practical benefits, particularly when they yield substantial performance improvements during large-scale adaptation. In contrast, the Clotho dataset serves as a complementary benchmark without any data overlap. For AC evaluation, we report a comprehensive suite of widely adopted captioning metrics, including BLEU (BLEU-1 to BLEU-4), FENSE, METEOR, ROUGE-L, CIDEr-D, SPICE, and SPIDEr, using the official evaluation toolkit [^2] https://github.com/Labbeti/aac-metrics.
@@ 3.3 Results and Analysis
To mitigate the potential influence of data overlap from pretraining and better reflect a realistic training pipeline—where a multimodal model is first pretrained on AC task before being further adapted to finer-grained tasks such as ADC or ACC—we fine-tuned the models on both the AudioCaps corpus (for AC) and one of the finer-grained datasets (ADC or ACC), as shown in Table 2 bottom. This setup enables the model to jointly learn general audio semantics and finer-grained distinctions.
From Table 2, we observe consistently stronger performance when applying ACC for fine-grained audio-text optimization alongside the traditional AC task, compared to the AC+ADC setting. This pattern holds across both the AudioCaps and Clotho datasets. On AudioCaps, ACC achieves substantially higher scores in CIDEr-D (2.9698 vs. 1.0475), SPIDEr (1.6904 vs. 0.6434), and BLEU metrics (e.g., BLEU-4: 0.2941 vs. 0.0976). On Clotho, although the differences are more modest, ACC still surpasses AC+ADC on key metrics such as BLEU-4 (0.0314 vs. 0.0271) and CIDEr-D (0.4034 vs. 0.3839). These results demonstrate that incorporating ACC as a fine-grained refinement task alongside AC consistently leads to more effective audio-text alignment and richer semantic understanding across diverse datasets, establishing ACC as a superior fine-tuning strategy for enhancing audio-text cross modal understanding in the context of multimodal large language models.
@@ 4 Conclusion
We propose Audio Commonality Captioning, a novel training paradigm for audio captioning that maintains fine-grained supervision while providing a gentler and more consistent training objective aligned with AC pretraining. Through comprehensive experiments on both in-domain and out-of-domain benchmarks, we show that ACC consistently outperforms standard Audio Captioning and Audio Difference Captioning in enhancing audio-text cross-modal understanding. Further evaluations demonstrate that ACC better preserves the model’s general capabilities across various speech and music-related tasks—achieving a more favorable balance between generalization and task-specific performance. These findings establish ACC as a effective and versatile training objective for advancing audio-text understanding in multimodal large language models.

### v3

@@ 2.1 Construction of the ADC and ACC Dataset
Qwen2-Audio + AC 0.1762 0.0894 0.0506 0.0295 0.5922 0.1529 0.2831 0.4132 0.1005 0.1762
Qwen2-Audio + ADC 0.0798 0.0262 0.0118 0.0053 0.3861 0.0566 0.0754 0.0943 0.0442 0.0798
Qwen2-Audio + ACC (ours) 0.1999 0.1039 0.0585 0.0338 0.5804 0.1534 0.2826 0.4118 0.1032 0.1999
@@ 3.1 Dataset and Metrics
In this study, we investigate the relative impact of direct Audio Captioning (AC) fine-tuning, Audio Difference Captioning (ADC), and Audio Commonality Captioning (ACC) on enhancing the audio-text cross-modal understanding of MLLMs. We adopt the audio captioning task as the primary evaluation benchmark, as it directly reflects the model’s ability to achieve semantic alignment between audio and text before and after fine-tuning. For evaluation, we utilize two widely used datasets AudioCaps [12] and Clotho [5]. It is worth noting that AudioCaps is partially incorporated into the construction of our fine-tuning datasets for AC, ADC, and ACC through sampling and mixing, which may introduce some overlap between pretraining and evaluation data. Nevertheless, within the context of MLLMs, such strategies remain valuable, as they offer meaningful empirical insights and practical benefits, particularly when they yield substantial performance improvements during large-scale adaptation. In contrast, the Clotho dataset serves as a complementary benchmark without any data overlap. For AC evaluation, we report a comprehensive suite of widely adopted captioning metrics, including BLEU (BLEU-1 to BLEU-4), FENSE, METEOR, ROUGE-L, CIDEr-D, SPICE, and SPIDEr, using the official evaluation toolkit [^2] https://github.com/Labbeti/aac-metrics.
@@ 3.3 Results and Analysis
To mitigate the potential influence of data overlap from pretraining and better reflect a realistic training pipeline, where a multimodal model is first pretrained on AC task before being further adapted to finer-grained tasks such as ADC or ACC, we fine-tuned the models on both the AudioCaps corpus (for AC) and one of the finer-grained datasets (ADC or ACC), as shown in Table 2 bottom. This setup enables the model to jointly learn general audio semantics and finer-grained distinctions. From Table 2, we observe consistently stronger performance when applying ACC for fine-grained audio-text optimization alongside the traditional AC task, compared to the AC+ADC setting. This pattern holds across both the AudioCaps and Clotho datasets. On AudioCaps, ACC achieves substantially higher scores in CIDEr-D (2.9698 vs.1.0475), SPIDEr (1.6904 vs.0.6434), and BLEU metrics (e.g., BLEU-4: 0.2941 vs.0.0976). On Clotho, although the differences are more modest, ACC still surpasses AC+ADC on key metrics such as BLEU-4 (0.0314 vs. 0.0271) and CIDEr-D (0.4034 vs. 0.3839). These results demonstrate that incorporating ACC as a fine-grained refinement task alongside AC consistently leads to more effective audio-text alignment and richer semantic understanding across diverse datasets, establishing ACC as a superior fine-tuning strategy for enhancing audio-text cross modal understanding in the context of multimodal large language models.
@@ 4 Conclusion
We propose Audio Commonality Captioning, a novel training paradigm for audio captioning that maintains fine-grained supervision while providing a gentler and more consistent training objective aligned with AC pretraining. Through comprehensive experiments on both in-domain and out-of-domain benchmarks, we show that ACC consistently outperforms standard Audio Captioning and Audio Difference Captioning in enhancing audio-text cross-modal understanding. Further evaluations demonstrate that ACC better preserves the model’s general capabilities across various speech and music-related tasks, achieving a more favorable balance between generalization and task-specific performance. These findings establish ACC as a effective and versatile training objective for advancing audio-text understanding in multimodal large language models.